# Techniques for improved statistical convergence in quantification of eddy diffusivity moments

### v1

@@ VI.7 Impact on constructing closure operators
where $a^{mn}(x_{2},t)$ are model coefficients determined using a process detailed in Liu et al. [8] and Lavacot et al. [9]. We use this form instead of truncating the Kramers–Moyal expansion in Equation (61) for several reasons. First, it is well known in the literature that the Kramers–Moyal expansion does not converge. That is, finite truncation of the expansion leads to divergent results. This property of the Kramers–Moyal expansion was proven by Pawula [24] and was shown to be true in modeling the eddy diffusivity by Liu et al. [8] and specifically in 2D RT instability by Lavacot et al. [9]. Secondly, a model involving truncation of the Kramers–Moyal expansion would be challenging to implement numerically in this spatiotemporal problem, since this would require time advancing spatial gradients of mixed derivatives. The inverse operator in Equation (67) addresses both of these issues. Once the coefficient fields $a^{mn}$ are determined using the MMI procedure, Equation (67) can be directly time-integrated using a partial differential equation solver. Details on this and the MMI process can be found in Liu et al. [8].
@@
In Lavacot et al. [9], the importance of eddy diffusivity moments for modeling mean scalar transport was investigated using the standard MFM. Different combinations of moments (i.e., different truncations of terms in the MMI operator in Equation (67)) were tested to assess the importance of each moment. Here, we examine the truncation of the MMI operator to the four terms shown in Equation (67). Construction of this model form requires the eddy diffusivity moments $D^{00}$, $D^{10}$, $D^{01}$, and $D^{20}$. We compare results of models constructed directly using measurements of $D^{mn}$ from the standard MFM and the decomposition MFM. For both methods, we use $D^{mn}$ measurements averaged over $200$ realizations. At this number of realizations, the moments are visually converged for the decomposition MFM, but not the standard MFM.
@@
Figure 8 shows mean concentration and turbulent scalar flux profiles resulting from each of the models. The results from the model using the standard MFM eddy diffusivity moments diverge significantly from the DNS results. This may lead to the incorrect conclusion that addition of terms in the MMI operator does not lead to convergence in this case. However, the source of this error is the large amount of statistical error in the higher-order moments. On the other hand, the results from the model using the decomposition MFM moments agree much better with the DNS. There appears to be some statistical error still at this number of realizations, but there is not nearly as much as there is with the standard MFM-based model. This highlights the importance of statistically-converged higher order moments in modeling. Compared to the standard MFM, the decomposition MFM is a more efficient method for obtaining statistically-converged moments that are usable for constructing models.
@@
Figure 8: (a) Mean concentration profiles and (b) turbulent scalar flux from DNS (solid black) and models using eddy diffusivity moments measured with the standard MFM (dashed green) and decomposition MFM (dash-dotted red). Results from both methods are averaged over $200$ realizations.
@@ VII Conclusion
In this work, modifications to the standard MFM are presented for faster statistical convergence of the eddy diffusivity moments. We first demonstrate the utility of using a single donor simulation for the receiver equations. Using a single donor prevents pileup of statistical error that may arise in separate donors, even when the same numerical initial conditions are used. Due to the potentially slow statistical convergence of MFM using separate donors, we recommend MFM using only one donor, though that may not be the natural implementation in some codes.
We also present formulations of the receiver equations based on the Kramers–Moyal expansion of the fluctuation field for both scalar and momentum transport—this is the decomposition MFM. Instead of solving for the scalar $c$, as in the standard MFM, the decomposition MFM solves for a variable based on the scalar fluctuation $\mathscr{c}$, allowing the mean forcing in the receiver equations to be handled semi-analytically. There is no additional cost to use decomposition MFM in place of the standard MFM, and the former has the added benefit that it also allows for consistent boundary condition treatment, such as in problems with periodic domains, like turbulent channel flow.
@@
To demonstrate its utility, we apply the decomposition and single donor MFM to 2D RT instability. In this case study, we demonstrate computational savings of at least an order of magnitude in reaching statistical convergence of eddy diffusivity moments compared to MFM using separate donors. We show that this improved statistical convergence is substantial and significantly impacts analysis of closure operators, as poorly-converged eddy diffusivity moments can lead to incorrect conclusions.
Acknowledgements. This work was performed under the auspices of the US Department of Energy by Lawrence Livermore National Laboratory under Contract No. DE-AC52-07NA27344. D.L. was additionally supported by the Charles H. Kruger Stanford Graduate Fellowship.
@@ Appendix A Error analysis
We present the following error analysis to illustrate the propagation of error due to multiple donor simulations. This analysis is done in one dimension ($x_{1}$) for simplicity. First, we define $F^{i}$ to be the measurements of the turbulent scalar flux used to determine $D^{i}$:
@@
In the following analysis, we use the notation $\varepsilon^{ij}$, where $i$ denotes the error associated with $\mathscr{c}^{i}$, as in Equation (71), and $j$ denotes the simulation used to determine $D^{j}$. The addition of index $j$ is introduced, because when different donor simulations are used in MFM, each simulation has its own statistical error. For example, $\varepsilon^{00}$ is not necessarily the same as $\varepsilon^{01}$, despite them both being errors associated with $\mathscr{c}^{0}$. The measurement of $D^{0}$ can then be written as
@@
If one donor simulation is used for all receiver equations in MFM, the third term above disappears, since $\varepsilon^{00}=\varepsilon^{01}=\varepsilon^{0}$. When different donors are used, $\varepsilon^{00}$ and $\varepsilon^{01}$ are not the same, so the overall statistical error scales with $x_{1}$, making statistical convergence for $D^{1}$ slower than for $D^{0}$. While this analysis is presented in one dimension, the variable $x_{1}$, can be considered as either a spatial or temporal variable. Since time can become large in numerical simulations, and measurements are often taken in late time, the statistical convergence of the first temporal moment is especially slow.
@@
If one donor is used, the fourth, fifth, and sixth terms vanish. If different donors are used, those terms remain and the overall statistical error scales by $x_{1}^{2}$, resulting in even slower statistical convergence for $D^{2}$.
@@
Appendix B Rayleigh–Taylor spatial eddy diffusivity moments
@@
Figure 9: Convergence of $D^{10}$ (normalized by $h$) MFM measurement using standard and decomposition methods for RT instability case.
@@
Figure 10: Convergence of $D^{20}$ (normalized by $h^{2}$) MFM measurement using standard and decomposition methods for RT instability case.

### v2

@@ VI.7 Impact on constructing closure operators
where $a^{mn}(x_{2},t)$ are model coefficients determined using a process detailed in Liu et al. [8] and Lavacot et al. [9]. We use this form instead of truncating the Kramers–Moyal expansion in Equation (61) for several reasons. First, it is well known in the literature that the Kramers–Moyal expansion does not converge. That is, finite truncation of the expansion leads to divergent results. This property of the Kramers–Moyal expansion was proven by Pawula [25] and was shown to be true in modeling the eddy diffusivity by Liu et al. [8] and specifically in 2D RT instability by Lavacot et al. [9]. Secondly, a model involving truncation of the Kramers–Moyal expansion would be challenging to implement numerically in this spatiotemporal problem, since this would require time advancing spatial gradients of mixed derivatives. The inverse operator in Equation (67) addresses both of these issues. Once the coefficient fields $a^{mn}$ are determined using the MMI procedure, Equation (67) can be directly time-integrated using a partial differential equation solver. Details on this and the MMI procedure can be found in Liu et al. [8].
@@
In Lavacot et al. [9], the importance of eddy diffusivity moments for modeling mean scalar transport was investigated using the standard MFM with separate donors. Different combinations of moments (i.e., different truncations of terms in the MMI operator in Equation (67)) were tested to assess the importance of each moment. Here, we examine the truncation of the MMI operator to the four terms shown in Equation (67). Construction of this model form requires the eddy diffusivity moments $D^{00}$, $D^{10}$, $D^{01}$, and $D^{20}$. We compare results of models constructed directly using measurements of $D^{mn}$ from the standard MFM with separate donors and the decomposition MFM. For both methods, we use $D^{mn}$ measurements averaged over $200$ realizations. At this number of realizations, the moments are visually converged for the decomposition MFM, but not the standard MFM using separate donors.
@@
Figure 8 shows mean concentration and turbulent scalar flux profiles resulting from each of the models. The results from the model using the separate-donor MFM eddy diffusivity moments diverge significantly from the DNS results. This may lead to the incorrect conclusion that addition of terms in the MMI operator does not lead to convergence in this case. However, the source of this error is the large amount of statistical error in the higher-order moments. On the other hand, the results from the model using the decomposition MFM moments agree much better with the DNS. There appears to be some statistical error still at this number of realizations but significantly less than with the MFM with separate donors. This highlights the importance of statistically-converged higher order moments in modeling. Compared to the standard MFM using separate donors, the decomposition MFM is a more efficient method for obtaining statistically-converged moments that are usable for constructing models.
@@
Figure 8: (a) Mean concentration profiles and (b) turbulent scalar flux from DNS (solid black) and models using eddy diffusivity moments measured using the standard MFM with separate donors (dashed green) and decomposition MFM (dash-dotted red). Eddy diffusivity moments from both methods are averaged over $200$ realizations.
@@ VII Conclusion
In this work, modifications to the standard MFM are presented for enhanced statistical convergence. We first demonstrate the utility of using a single donor simulation for the receiver equations. Using a single donor prevents pileup of statistical error that may arise in separate donors, even when the same numerical initial conditions are used. Due to the potentially slow statistical convergence of MFM using separate donors, we recommend MFM using only one donor, though that may not be the natural implementation in some codes. In cases where the repeated numerical solutions of the donor equations produce identical instantaneous results (see Appendix B), using a single donor is not strictly necessary; however, it does have the added cost-saving benefit of not solving the donor equations multiple times. Moreover, for cases where the MFM forcing may be incompatible with the boundary conditions of the problem, decomposition MFM must be used which would also require using a single donor.
While the statistical error issue presented in this work is due to computational processes specific to the Ares code, sources of error between separate donors are not limited to computation. For example, when performing MFM for chaotic flows requiring multiple realizations for statistical convergence, one may consider computing new higher-order moments using a new donor but keeping previously-computed lower-order moments from a separate donor. In this case, there are certainly differences between the donors of the lower-order and higher-order moments. As we have shown in this work, small differences between separate donors significantly affect statistical error, and one should expect to incur larger statistical error in the higher-order moments. Another interpretation of our results is that since differences in ensembles can result in amplified statistical error in the moments, simulations belonging to one ensemble should remain with that ensemble, i.e., they should not be used for statistics in a different ensemble. Overall, this work highlights the nuances with the donor equations that one should consider to avoid additional statistical error when using MFM to compute the eddy diffusivity moments.
We also present decomposition MFM, which naturally uses a single donor, for both scalar and momentum transport. For example, for scalar transport, instead of solving for the scalar $c$ as in standard MFM, decomposition MFM solves for a variable based on the scalar fluctuation $\mathscr{c}$, which is then expanded using the Kramers-Moyal expansion, allowing the mean forcing in the receiver equations to be handled semi-analytically. Decomposition MFM allows for consistent boundary condition treatment, such as in problems with periodic domains.
@@
To demonstrate its utility, we apply decomposition MFM to 2D RT instability. In this case study, we demonstrate computational savings of at least an order of magnitude in reaching statistical convergence of eddy diffusivity moments compared to MFM using separate donors. This improved statistical convergence is since decomposition MFM inherently uses a single donor rather than separate donors. We show that this improved statistical convergence is substantial and significantly impacts analysis of closure operators, as poorly-converged eddy diffusivity moments can lead to incorrect conclusions.
Acknowledgements. This work was performed under the auspices of the US Department of Energy by Lawrence Livermore National Laboratory under Contract No. DE-AC52-07NA27344. D.L. was additionally supported by the Charles H. Kruger Stanford Graduate Fellowship. We are also thankful to Dr. Sho Takatori for discussions on MFM and pointing us to connections to Generalized Taylor Dispersion Theory in the work of Dr. Howard Brenner.
@@ Appendix A Error analysis
We present the following error analysis to illustrate the propagation of error due to separate donor simulations. This analysis is done in one dimension ($x_{1}$) for simplicity. First, we define $F^{i}$ to be the measurements of the turbulent scalar flux used to determine $D^{i}$:
@@
In the following analysis, we use the notation $\varepsilon^{ij}$, where $i$ denotes the error associated with $\mathscr{c}^{i}$, as in Equation (71), and $j$ denotes the simulation used to determine $D^{j}$. The addition of index $j$ is introduced, because when separate donor simulations are used in MFM, each simulation has its own statistical error. For example, $\varepsilon^{00}$ is not necessarily the same as $\varepsilon^{01}$, despite them both being errors associated with $\mathscr{c}^{0}$. The measurement of $D^{0}$ can then be written as
@@
If one donor simulation is used for all receiver equations in MFM, the third term above disappears, since $\varepsilon^{00}=\varepsilon^{01}=\varepsilon^{0}$. When separate donors are used, $\varepsilon^{00}$ and $\varepsilon^{01}$ are not the same, so the overall statistical error scales with $x_{1}$, making statistical convergence for $D^{1}$ slower than for $D^{0}$. While this analysis is presented in one dimension, the variable $x_{1}$, can be considered as either a spatial or temporal variable. Since time can become large in numerical simulations, and measurements are often taken in late time, the statistical convergence of the first temporal moment is especially slow.
@@
If one donor is used, the fourth, fifth, and sixth terms vanish. If separate donors are used, those terms remain and the overall statistical error scales by $x_{1}^{2}$, resulting in even slower statistical convergence for $D^{2}$.
@@
Appendix B Turbulent channel flow case study: Effect of using separate donors vs. a single donor and compatibility with decomposition MFM
Figure 9: Comparison of using standard MFM with separate donors vs. a single donor for quantifying the first and second spatial moments of the nonlocal eddy viscosity in turbulent channel flow at $Re_{\tau}=180$. The converged solution using standard MFM with a single donor at $1000$ eddy turnover time ($1000T$) is shown for comparison. Using separate donors clearly has more error at $100T$ than using a single donor.
Figure 10: Comparison of using standard MFM with a single donor vs. decomposition MFM for quantifying the first spatial moment of the nonlocal eddy viscosity in turbulent channel flow at $Re_{\tau}=180$. Standard MFM with a single donor and decomposition MFM are almost identical at $100T$.
In this appendix, we demonstrate the effect of using separate donors vs. a single donor in the context of turbulent channel flow. We then examine the effect of using decomposition MFM in conjuction with the single donor simulations. This test case also serves as a demonstration of decomposition MFM for momentum as developed in Section V.
The turbulent channel flow does not suffer from the processor issues seen in the RT case, i.e., repeated numerical solutions of the turbulent channel flow donor equations with the same initial condition are identical and using separate donors would produce the same result as using a single donor. For the purposes of evaluating the equivalent effect of “separate donors” vs. “single donor,” we introduce differences in the initial condition. “Separate donors” uses different initial conditions for the donor simulations whereas “single donor” uses the same initial condition for all donor simulations. While truly using a single donor, e.g., the implementation shown in Figure 1b, is not strictly necessary for this case, it does have the cost-saving advantages of not repeatedly solving the donor equations. Moreover, using a single donor enables the use of decomposition MFM which is needed for quantifying eddy viscosity components related to the streamwise gradient of the mean velocity for which the periodic boundary conditions are incompatible with the MFM forcing [11, 13].
Similar to the setup of Park and Mani [11], we use a friction Reynolds number of $Re_{\tau}=180$ where $\tau=u_{\tau}/h$ and where $u_{\tau}$ is the friction velocity and $h$ is the channel half-height. Averaging is taken over the homogeneous streamwise ($x_{1}$) and spanwise ($x_{3}$) directions, and the Reynolds stresses are only a function of the wall-normal ($x_{2}$) direction. Figure 9 shows a comparison of standard MFM using separate donors vs. a single donor for the first and second moments of the eddy viscosity in the wall-normal direction, $D^{10}_{2121}$ and $D^{20}_{2121}$, respectively. The tensor subscripts correspond to the shear component of the Reynolds stress, $\langle u_{2}^{\prime}v_{1}^{\prime}\rangle$, and the velocity gradient, $\partial\langle v_{1}\rangle/\partial x_{2}$, as shown in Equation (39). The solution at $1000T$ is taken to be the converged solution where $T$ is the eddy turnover time. Figure 9 clearly shows the amplified errors associated with using separate donors.
Figure 10 shows a comparison between standard MFM using a single donor and decomposition MFM. The momentum equations for decomposition MFM are detailed in Equations (45a)-(46b). The two methods have almost identical solutions at $100T$, which is expected since standard MFM and decomposition MFM are mathematically equivalent. All improvement in statistical convergence is due to the usage of a single donor rather than separate donors.
Appendix C Rayleigh–Taylor spatial eddy diffusivity moments
@@
Figure 11: Convergence of $D^{10}$ (normalized by $h$) measurement using standard MFM with separate donors and decomposition MFM for the RT instability case.
@@
Figure 12: Convergence of $D^{20}$ (normalized by $h^{2}$) measurement using standard MFM with separate donors and decomposition MFM for the RT instability case.